The moment of inertia of a square lamina about the perpendicular axis through its centre of mass is 20kg then it`s moment of inertia about an axis touching it`s side and in the plane of lamina will be

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
Given parameters

Moment of inertia of a square lamina about the perpendicular axis at the center of the square (I0) = 20 kgm-2

Let us consider

L = Length of square

M = Mass of the square laminaI0=ML2620 = ML2/6

ML2= 120 kgm2

According to the parallel axis theorem

Ix= Iy= I0/2= 10kgm-2

I1= Ix+ M(L/2)2

⇒ I1= Ix+ ML2/4

⇒ I1= 10 + 120/4

⇒ I1= 40 kgm-2